# autocorrelation function CFR channel frequency response

### v3

@@ Abstract
This paper focuses on the parameterization of the multipath propagation model (MPM) for indoor broadband power line communications (PLC), which up to now has been established in an heuristic way. The MPM model was originally proposed for outdoor channels in the band up to 20 MHz, but its number of parameters becomes extremely large when used to model indoor channel frequency responses (CFR), which are much more frequency-selective than outdoor ones, and the band is extended to 80 MHz. This work proposes a fitting procedure that addresses this problem. It allows determining the model parameters that yield the best fit to each channel of a large database of single-input single-output (SISO) experimental measurements acquired in typical home premises of different European countries. Then, the properties of the MPM parameters are analyzed. The study unveils the relation between the model parameters and the main characteristics of the actual CFR like the frequency selectivity and the average attenuation. It also estimates the probability density function (PDF) of each parameter and proposes a fitting distribution for each of them. Moreover, the relationship among the main parameters of the model is also explored. Provided results can be helpful for the development of MPM-based models for indoor broadband PLC.
Index Terms:
Power line communications (PLC), channel model, multipath propagation model, channel frequency response, statistical characterization, probability density function.
@@ I Introduction
The increasing demand of high-speed multimedia services led to reconsider the indoor power delivery network as a data transmission medium. The power line communication (PLC) technology is becoming an established solution since it exploits the existing wired infrastructure to convey high-speed data content, thus saving costs and deployment time. Furthermore, the latest standards and devices include advanced signal processing and enhanced transmission techniques, such as the exploitation of all the deployed network conductors and the bandwidth extension, that improve even further the achievable data rate, guaranteeing transmission rates in the order of gigabit per second at the physical layer [1, 2, 3]. This high performance level motivates a renewed interest in developing novel and updated PLC solutions, especially for home networking and smart grid applications [4, 5, 6, 7, 8].
@@
In order to support the growth of PLC technology, it is fundamental to provide effective channel models, which are able to numerically emulate the properties of actual PLC networks avoiding on field tests. Two basic approaches can be tackled in defining a model, i.e., bottom-up and top-down, and both of them can follow a deterministic or a statistical strategy [9, 10, 11, 12]. A deterministic strategy typically considers a specific network configuration, while a statistical strategy attempts to introduce some variability by means of statistical distributions. Concerning the possible approaches, the bottom-up method has a strong physical connection and describes the phenomena involved in the signal propagation by exploiting the transmission line theory. Typically, it provides a faithful channel representation, although it turns out to be a computationally onerous procedure. Some examples concerning the single-input-single-output (SISO) channel modeling can be found in [13, 14], which propose a deterministic approach, and in [15, 16, 17, 18, 11], which follow a statistical approach. Recently, the models discussed in [19] and [20] extends the SISO bottom-up strategy to the multiple-input multiple-output (MIMO) case.
@@
Conversely, the top-down approach is an empirical method that fits an analytic function to a measurements database. This strategy facilitates the statistical extension by adopting a certain distribution for the parameters that describe the theoretical function. One of the first and more common SISO models is the multipath propagation model (MPM) proposed in [9] and was later statistically extended in [21]. Extensions of the SISO formulation to the MIMO are presented in [22] and [23].
@@
The PLC channel response also exhibits a short-term variation synchronous with the mains voltage, which allows modelling it as a linear periodically time-varying (LPTV) system [24]. However, since the time variation is quite small compated to the duration of the impulse response, i.e., the channel response is underspread, the slow variation approximation can be made and the channel response can be modeled as a series of time-invariant responses that repeat with the period of the mains signal. Hence, the MPM model can be applied to each of these time-invariant responses.
@@ I-A Contribution
The top-down MPM model formulated in [9] has been widely used in the PLC scenario [25, 10, 26, 27, 28]. However, it was proposed for outdoor PLC channels in the band up to $20$ MHz. Its statistical extension to indoor channels made in [29] and [21] was based on heuristic approaches and no rigorous measurement-based analysis of this model for indoor scenarios in the band up to 80 MHz has been found in the literature.
@@
  • We propose a method to determine the parameters of the MPM model for indoor broadband PLC channels in the band up to $80$ MHz. It addresses the problem of the extremely large number of parameters that results when the MPM model in [9] is extended to indoor scenarios and that are a consequence of the much more frequency selective behavior of indoor channel frequency responses and of the bandwidth extension to $80$ MHz. Since interference with commercial frequency modulation (FM) broadcasting is important above $87$ MHz in many regions.
@@
  • We give an in-depth analysis of the resulting MPM parameters, showing their relation to the main channel characteristics like the frequency selectivity and the average attenuation, assessing their statistical behavior and the relation between them. Unlike the a posteriori parameters inference discussed in previous works, this study unveils the intrinsic properties of the MPM parameter and can be helpful for the development of novel MPM models or the refinement of existent ones.
@@ I-B Organization
The paper is organized as follows. First, the MPM is briefly recalled in Section II. Second, the assumptions and the method proposed to determine the MPM parameters that give the best fit to each measured channel is detailed and justified in Section III. The properties of the computed parameters, such as their relation to the channel characteristics, statistical behavior and relation between them are discussed in Section IV. Finally, conclusions follow in Section V.
@@ II Background: The Multipath Propagation Model
The MPM follows a top-down approach and describes the multipath propagation of the power signal into the PLC medium that is due to the line discontinuities, such as branches, or unmatched loads. The analytic formulation of the CFR can be expressed as [9]
@@
where $a_{0}$ and $a_{1}$ are the attenuation coefficients, $g_{i}$ and $d_{i}$ are the path gain and the path length of the $i$-th path [^1] For the sake of notation simplicity in the subsequent fitting process, paths have been indexed as $10\leq i\leq N-$ 1, instead of $1\leq i\leq N$ as done in [9]., respectively. The coefficient $g_{i}$ can be interpreted as the product of transmission and reflection coefficients by restricting $|g_{i}|\leq 1$ and by introducing the normalization coefficient $A>0$, which allows adjusting the attenuation of the resulting CFR [9]. The term $N$ denotes the number of paths and $K$ is the exponent of the attenuation factor related to the cable features. This work uses $N=2554$ (its rationale will be justified in Section III.III-B), which is much larger than the values used in outdoor channels. Typical values of $K$ range between 0.5 and 1 [9], hence $K=1$ is assumed for the rest of this paper. The parameter $\nu=c\epsilon_{r}$ is the propagation speed of light in the cables, where $c$ is the speed of light in the vacuum and $\epsilon_{r}$ is the relative dielectric constant of the insulator which envelops the conductors, which has been set to $\epsilon_{r}=1.5$ in order to take into account the non-uniformity of the dielectric [9].
III Determination of the MPM Parameters
@@
In order to provide a statistical characterization of the parameters that describe the MPM formulation in (1), the first step is to compute the MPM parameters that give the best fitting of this analytic expression to each of the channels of the measurement database. The latter consists of $426$ indoor CFRs measured in different European countries using a vector network analyzer. CFRs have $M=1262$ frequency samples given by $f_{m}=f_{0}+m\Delta f$, with $f_{0}=1.0$ MHz, $10\leq m\leq M-$ 1 and $\Delta f=62.5978$ kHz. Part of this collection was previously used in [23] and [12], and the results are consistent with the ones obtained from other extensive measurements in Italy that were characterized in [30].
@@
Since the model is described by many parameters, the following assumptions must be adopted in order to simplify the fitting strategy. First, the attenuation coefficients $a_{0}$ and $a_{1}$ are estimated through the robust regression fit of each channel measurement. This linear attenuation profile can assumed to be essentially determined by the longest path of the MPM. Hence, we consider the model in (1) with a single path of the largest length, $L$ (its value will be given in Section III.III-B). Since $A$ is a normalization factor used to force $|g_{i}|\leq 1$, when a single path is considered it happens that $|A\cdot g_{i}|=1$. Accordingly, the robust regression fit of the longest path in (1) can be expressed as
@@
where $\alpha_{0}$ and $\alpha_{1}$ are the coefficients of the robust regression fit. Thus, the attenuation coefficients are computed as
@@
This is a viable assumption since indoor PLC networks typically exhibit a complex topology with a great number of branches. This results into approximately uniformly distributed distances (spaced $1.25$ m for the values of $L$ and $N$ considered in this work) traveled by the signal and its reflections from the transmitter side towards the receiver outlet.
Afterwards, the path gains, $g_{i}$ that give the best fitting to each each measured channel have to be determined. This could be done by means of a LS estimation that minimizes the root mean square error (RMSE) w.r.t. the measured channel in the set of $M$ measured frequencies $f_{m}$,
@@
The $g_{i}$ can be obtained by solving a weighted (WLS) problem formulated by expressing (1) in matrix form as follows,
@@ III-A Dominant Paths Selection Procedure
The rationale for this criterion is that the actual contribution of a path to the overall CFR is essentially determined by the factor $|g_{i}|\cdot\exp(-(a_{0}+a_{1}f)d_{i}$ in the MPM in (1).
The decimation procedure continues while the NRMSE (dB) remains lower than a given threshold, which has been set to $-20$ dB. This ensures that the channel computed with a reduced set of paths still represents a very good approximation to the corresponding measured channel. Since the columns of the matrix $\mathbf{P}$ are not orthogonal, the path gains $g_{i}$ have to be recomputed after a path is discarded. To this end, the problem in (7) is reformulated by removing the $j$-th column of the matrix $\mathbf{P}$ and the new path gains are obtained. Table I summarizes the presented fitting and decimation procedure applied to each measured channel, where the expressions to obtain the values of $L$ and $N$ will be given in Section III.III-B.
@@
7: while NRMSE (dB)$<\text{NRMSE}_{\text{threshold}}$ do
@@
The number of initial paths considered for the model at the pre-decimation point is $N=2554$ (as justified in the next subsection). This gives average and standard deviation values of the NRMSE (computed over the set of measured channels) equal to $\overline{\textrm{NRMSE}}=-36.79$ dB and $\sigma_{\textrm{NRMSE}}=7.99$ dB, respectively. After the application of the proposed decimation procedure, the number of paths is reduced to an average value (over the whole set of fitted channels) of 217.84, i.e., less than 10% of the initial value, and the corresponding average and standard deviation values of the NRMSE (dB) are $\overline{\textrm{NRMSE}}=-20.05$ dB and $\sigma_{\textrm{NRMSE}}=0.0678$ dB, respectively. Hence, the proposed decimation process achieves a good trade-off between modeling complexity and accuracy.
For illustrative purposes, Fig.1 shows the amplitude and unwrapped phase of a measured CFR, $H(f)$, and its corresponding fitted response, $\hat{H}(f)$, obtained with $N=183$ dominant paths determined using the proposed decimation process. As can be observed, the measured trace is almost veiled by the estimated one, which highlights the notable accuracy of the fitting process.

### v4

# autocorrelation function AWGN additive white Gaussian noise

@@ Abstract
This paper focuses on the parameterization of the multipath propagation model (MPM) for indoor broadband power line communications (PLC), which up to now has been established in an heuristic way. The MPM model was initially proposed in the PLC context for outdoor channels in the band up to 20 MHz, but its number of parameters becomes extremely large when used to model indoor channel frequency responses (CFR), which are much more frequency-selective than outdoor ones, and the band is extended to 80 MHz. This work proposes a fitting procedure that addresses this problem. It allows determining the model parameters that yield the best fit to each channel of a large database of single-input single-output (SISO) experimental measurements acquired in typical home premises of different European countries. Then, the statistics of the MPM parameters are analyzed. The study unveils the relation between the model parameters and the main characteristics of the actual CFR like the frequency selectivity and the average attenuation. It also estimates the probability density function (PDF) of each parameter and proposes a fitting distribution for each of them. Moreover, the relationship among the main parameters of the model, as well as their impact on the performance of PLC communication systems are also explored. Provided results can be helpful for the development of MPM-based models for indoor broadband PLC.
@@ I Introduction
The increasing demand of high-speed multimedia services led to reconsider the indoor power delivery network as a data transmission medium. The power line communication (PLC) technology is becoming an established solution since it exploits the existing wired infrastructure to convey high-speed data content, thus saving costs and deployment time. Furthermore, the latest standards and devices include advanced signal processing and enhanced transmission techniques, such as the exploitation of all the deployed network conductors and the bandwidth extension, that improve even further the achievable data rate, guaranteeing transmission rates in the order of gigabit per second at the physical layer [Yonge13, G.9960, IEEE_P1901]. This high performance level motivates a renewed interest in developing novel and updated PLC solutions, especially for home networking and smart grid applications [Bush2013, Prasad2020, Righini2021, Canete2023, Cortes23bis].
@@
In order to support the growth of PLC technology, it is fundamental to provide effective channel models, which are able to numerically emulate the properties of actual PLC networks avoiding on field tests. Two basic approaches can be tackled in defining a model, i.e., bottom-up and top-down, and both of them can follow a deterministic or a statistical strategy [Zimmermann02, Tonello07bis, Canete2011, Povedano23]. A deterministic strategy typically considers a specific network configuration, while a statistical strategy attempts to introduce some variability by means of statistical distributions. Concerning the possible approaches, the bottom-up method has a strong physical connection and describes the phenomena involved in the signal propagation by exploiting the transmission line theory. Typically, it provides a faithful channel representation, although it turns out to be a computationally onerous procedure. Some examples concerning the single-input-single-output (SISO) channel modeling can be found in [Canete02, Galli06], which propose a deterministic approach, and in [Philipps99, Philipps00, Esmailian03, Tonello11, Canete2011], which follow a statistical approach. The models discussed in [Versolatto11] and [Corchado16] extend the SISO bottom-up strategy to the multiple-input multiple-output (MIMO) case.
@@
Conversely, the top-down approach is an empirical method that fits an analytic function to a measurements database. This strategy facilitates the statistical extension by adopting a certain distribution for the parameters that describe the theoretical function. The multipath propagation model (MPM) is an example of this approach, and its validity in both outdoor and indoor PLC scenarios is supported by the fact that power line cables are electrically long at the considered frequencies, resulting in a received signal that consists of multiples echoes caused by impedance mismatches at multiple points in the network. One of the first and more common MPM SISO models was proposed in [Zimmermann02] and later statistically extended in [Tonello12]. Extensions of the SISO formulation to the MIMO case are presented in [Canova2010] and [Pagani16].
@@
The PLC channel response also exhibits a short-term variation synchronous with the mains voltage, which allows modelling it as a linear periodically time-varying (LPTV) system [Canete06]. However, since the time variation is quite small compared to the duration of the impulse response, i.e., the channel response is underspread, the slow variation approximation can be made and the channel response can be modeled as a series of time-invariant responses that repeat with the period of the mains signal. Hence, the MPM model can be applied to each of these time-invariant responses.
@@ I-A Contribution
The MPM model has been widely used in PLC scenarios [Zimmermann02, OperaChannel05, Tonello07bis, Chien15, Huang2012, Gay2016]. Its validity for outdoor channels in the frequency band up to $20$ MHz was empirically assessed in [Zimmermann02], which provided the model parameters that yielded the best fit to a set of measured channels. Statistical extensions to indoor channels were made in [Tonello07] and [Tonello12], which proposed channel generators that assumed certain a priori probability distributions for the model parameters based on heuristic approaches. However, a rigorous measurement-based analysis neither of the MPM capacity to model indoor scenarios in the band up to 80 MHz nor of the statistical distributions of the model parameters has been found in the literature.
@@
  • We propose a method to determine the parameters of the MPM model for indoor broadband PLC channels in the band up to $80$ MHz. It addresses the problem of the extremely large number of parameters that results when the MPM model in [Zimmermann02] is extended to indoor scenarios and that are a consequence of the much more frequency selective behavior of indoor channel frequency responses and of the bandwidth extension to $80$ MHz.
@@
  • We give an in-depth analysis of the resulting MPM parameters, showing their relation to the main channel characteristics like the frequency selectivity and the average attenuation, assessing their statistical behavior and the relation between them. Unlike the a posteriori parameters inference discussed in previous works, this study unveils the intrinsic properties of the MPM parameter and can be helpful for the development of novel MPM models or the refinement of existing ones. A discussion in this regard, and a study of the relationship between the main MPM parameters and the performance of indoor PLC systems are also provided.
@@ I-B Organization
The paper is organized as follows. First, the MPM is briefly recalled in Section II. Second, the assumptions and the method proposed to determine the MPM parameters that give the best fit to each measured channel is detailed and justified in Section III. The properties of the computed parameters, such as their relation to the channel characteristics, statistical behavior and relation between them are discussed in Section IV. The application of the proposed results to a further development of MPM-based channel generators and the analysis of the influence of the model parameters in the performance of PLC systems is discussed in Section V. Finally, conclusions follow in Section VI.
@@ II Background: The Multipath Propagation Model
The MPM follows a top-down approach and describes the multipath propagation of the power signal into the PLC medium that is due to the line discontinuities, such as branches, or unmatched loads. The analytic formulation of the CFR can be expressed as [Zimmermann02]
@@
where $a_{0}$ and $a_{1}$ are the attenuation coefficients, $g_{i}$ and $d_{i}$ are the path gain and the path length of the $i$-th path [^1] For the sake of notation simplicity in the subsequent fitting process, paths have been indexed as $0\leq i\leq N-1$, instead of $1\leq i\leq N$ as done in [Zimmermann02]., respectively. The coefficient $g_{i}$ can be interpreted as the product of transmission and reflection coefficients by restricting $|g_{i}|\leq 1$ and by introducing the normalization coefficient $A>0$, which allows adjusting the attenuation of the resulting CFR [Zimmermann02]. The term $N$ denotes the number of paths and $K$ is the exponent of the attenuation factor related to the cable features. This work uses $N=2554$ (its rationale will be justified in Section III.III-B), which is much larger than the values used in outdoor channels. Typical values of $K$ range between 0.5 and 1 [Zimmermann02], hence $K=1$ is assumed for the rest of this paper. The parameter $\nu=c\epsilon_{r}$ is the propagation speed of light in the cables, where $c$ is the speed of light in the vacuum and $\epsilon_{r}$ is the relative dielectric constant of the insulator which envelops the conductors, which has been set to $\epsilon_{r}=1.5$ in order to take into account the non-uniformity of the dielectric [Zimmermann02].
Expression (1) is a truncated version of the actually infinite number of forward- and backward-traveling waves caused by the reflections due to impedance mismatch. The rationale for this truncation is that the amplitude of these waveforms decreases at each reflection. Hence, those that have experienced a larger number of reflections would have negligible amplitude and can be ignored. It must be emphasized that the path lengths, $d_{i}$, correspond to the actual distances traveled by the referred waves. For instance, in a simple network consisting of a voltage source and a mismatched transmission line of length $\mathcal{L}$ terminated in a load impedance, the distances travelled by the received waves are $d_{i}=(2i+1)\mathcal{L}$, with $i=0,1,2,\ldots$ [Paul94, Ch. 5].
III Method to Determine the MPM Parameters
@@
In order to provide a statistical characterization of the parameters that describe the MPM formulation in (1), the first step is to compute the MPM parameters that give the best fitting of this analytic expression to each of the channels of the measurement database. It consists in a set of $426$ indoor CFRs corresponding to differential transmission and reception between the phase and neutral conductors (P-N), out of which $205$ have been acquired in Spain and $221$ where measured in the ETSI STF 410 MIMO PLC campaign described in [Pagani16]. Measured CFRs consists of $M=1262$ frequency samples given by $f_{m}=f_{0}+m\Delta f$, with $f_{0}=1.0$ MHz, $0\leq m\leq M-1$ and $\Delta f=62.5978$ kHz.
The characteristics of the measured channels are similar to those of the extensive campaign carried out in Italy [Tonello14] and to the ones measured in the United States and reported in [Galli11]. They also exhibit well-known properties of indoor PLC channels, such as the lognormal distribution of the delay spread, which has a mathematical foundation [Galli11]. The features of the employed channels are also coherent with the ones in [Pagani16]. However, it is worth mentioning that while the $351$ channels used in the latter to derive a MIMO channel model were also acquired in the referred ETSI campaign, they correspond to the P, N, E and CM (common mode) receiver ports described in [ETSI_TR_101_562-3], while in this work the conventional SISO differential P-N links are employed.
@@
Since the model is described by many parameters, the following procedure is used to simplify the fitting strategy. First, the attenuation coefficients $a_{0}$ and $a_{1}$ are estimated. To this end, we consider an MPM consisting on a single path. By setting $N=1$ in (1), it happens that $|A\cdot g_{i}|=1$, since $A$ is a normalization factor used to force $|g_{i}|\leq 1$. Furthermore, $N=1$ yields a linear attenuation profile from which the attenuation coefficients $a_{0}$ and $a_{1}$ can be estimated through the regression fit of each channel measurement. This linear attenuation profile can be assumed essentially determined by the longest path of the MPM, which is denoted by $d_{1}=L$ (its value will be given in Section III-B). The regression fit can be expressed as
@@
where $\alpha_{0}$ and $\alpha_{1}$ are the coefficients of the regression fit. Thus, the attenuation coefficients are computed as
@@
The presented procedure does not impose a rigid constraint on $a_{0}$, as there are multiple values of $a_{0}$, $g_{i}$ and $d_{i}$ that yield the same $g_{i}e^{-a_{0}d_{i}}$. Hence, if the longest path of the considered channel is shorter than $L$, the small value of $a_{0}$ resulting from the employed procedure can be compensated with larger values of $g_{i}$. Regarding $a_{1}$, which is related to the slope of the attenuation profile, its value is generally small, since the attenuation in indoor PLC channels does not exhibit the remarkable frequency dependent profile of outdoor channels [Cortes23, Fig.5], as involved distances are much shorter and attenuation is mainly due to multipath propagation and not to the skin effect. The validity of these assumptions will be corroborated later by the fitting results.
@@
It should be noted that actual path lengths are inherently discrete, as the lengths of the cable sections through which the waves propagate (forward and backward) are discrete quantities. Since indoor PLC networks have complex topologies with numerous branches whose lengths are unknown a priori, the uniform distribution appears to be the most appropriate choice. This assumption defines a grid of potential path lengths. Actual path lengths would be rounded to these grid elements. For the values of $L$ and $N$ used in this work (which will be given in Section III-B), the spacing is $1.25$ m. This means that the maximum rounding error of a path length would be $1.25/2=0.625$ m, which seems to be adequate even for small power line networks. Furthermore, it must be emphasized that the uniform grid is used as a starting point, those $d_{i}$ that do not correspond to actual (rounded) propagation distances within the considered network will have $g_{i}\approx 0$ and will be pruned in subsequent steps of the procedure.
Afterwards, the path gains, $g_{i}$, that give the best fitting to each measured channel have to be determined. This could be done by means of a LS estimation that minimizes the root mean square error (RMSE) w.r.t. the measured channel in the set of $M$ measured frequencies $f_{m}$,
@@
The gains $g_{i}$ can be obtained by solving a weighted (WLS) problem formulated by expressing (1) in matrix form as follows,
@@ III-A Dominant Paths Selection Procedure
The rationale for this criterion is that the contribution of a path to the overall CFR is determined by $|g_{i}|e^{-(a_{0}+a_{1}f)d_{i}}$ and not solely by the path length. Indeed, results presented in Section IV-B corroborate that the longest path is not generally the one with the smallest contribution.
The decimation procedure continues while the NRMSE (dB) remains lower than a given threshold, which has been set to 20-20 dB. This ensures that the channel computed with a reduced set of paths still represents a very good approximation to the corresponding measured channel. Since the columns of the matrix $\mathbf{P}$ are not orthogonal, the path gains $g_{i}$ have to be recomputed after a path is discarded. To this end, the problem in (7) is reformulated by removing the $j$-th column of the matrix $\mathbf{P}$ and the new path gains are obtained. Table I summarizes the presented fitting and decimation procedure applied to each measured channel, where the expressions to obtain the values of $L$ and $N$ will be given in Section III-B.
@@
[…]
@@
Table II summarizes the results obtained when the aforementioned decimation path procedure is applied to fit the measured channels in the frequency bands up to $20$ MHz and up to $80$ MHz. The former is included for comparison with the seminal work in [Zimmermann02], which applied the MPM to outdoor scenarios in the frequency band up to $20$ MHz. The number of initial paths at the pre-decimation stage is $N=640$ for the $20$ MHz bandwidth and $N=2554$ for the $80$ MHz (as justified in the next subsection). As expected, the average number of paths after the decimation process is much smaller in the $20$ MHz case (less than half) than in the $80$ MHz one. Also, its variance $\sigma_{N}$ is reduced. However, the number of paths is still much larger than in outdoor scenarios [Zimmermann02], where channels are less frequency selective because of the reduced number of derivations and increased attenuation (due to longer distances) that reduce the destructive effect of the multipath propagation.
Regarding the efficiency of the decimation procedure, the number of paths is reduced to an average value (over the whole set of fitted channels) of $107.97$ in the $20$ MHz case and to $217.84$ in the $80$ MHz bandwidth. Interestingly, the former is about 16.816.8% of the initial $N$, while the latter is lower than 1010% of the initial $N$. The higher efficiency in the $80$ MHz band is a consequence of the lower frequency selectivity of indoor PLC channels above 30 MHz [Cortes23bis].
[…]
The running time of the iterations of the decimation procedure [^3] Executed by MATLAB R2024b in a computer with Intel core i9-10900 CPU at 2.80 GHz and 32 GB of RAM. quadratically decreases from approximately $7$ s for the first one to less than $0.1$ s for iterations above $2000$, since the size of the matrices in (9) get smaller as the number of decimated paths increases. Nevertheless, it must be emphasized that the presented decimation process is not subject to any real-time constraint, as it might be the case of a real-time MPM channel generator.
Table II: Summary of the dominant path selection procedure for channels up to $20$ MHz and $80$ MHz
Frequency band
Magnitude 20 MHz 80 MHz
Value of $N$ before decimation 640 2554
$\overline{\textrm{NRMSE}}$ (dB) before decimation -30.77 -36.79
$\sigma_{\textrm{NRMSE}}$ (dB) before decimation 7.69 7.99
Average $N$ after decimation 107.97 217.84
$\sigma_{N}$ (dB) after decimation 55.50 134.58
$\overline{\textrm{NRMSE}}$ (dB) after decimation -20.11 -20.05
$\sigma_{\textrm{NRMSE}}$ (dB) after decimation 0.1537 0.0687
[…]
Presented results shows that the proposed decimation process achieves a good trade-off between modeling complexity and accuracy. For illustrative purposes, Fig. 2 shows the amplitude and unwrapped phase of a measured CFR, $H(f)$, and its corresponding fitted response, $\hat{H}(f)$, obtained with $N=183$ dominant paths determined using the proposed decimation process. As can be observed, the measured trace is almost veiled by the estimated one, which highlights the notable accuracy of the fitting process.